# Meaningful Learning: Advancing Abstract Reasoning in Large Language Models via Generic Fact Guidance

### v1

@@ Abstract
Large language models (LLMs) have developed impressive performance and strong explainability across various reasoning scenarios, marking a significant stride towards mimicking human-like intelligence. Despite this, when tasked with simple questions supported by a generic fact, LLMs often fail to provide consistent and precise answers, indicating a deficiency in abstract reasoning abilities. This has sparked a vigorous debate about whether LLMs are genuinely reasoning or merely memorizing. In light of this, we design a preliminary study to quantify and delve into the abstract reasoning abilities of existing LLMs. Our findings reveal a substantial discrepancy between their general reasoning and abstract reasoning performances. To relieve this problem, we tailor an abstract reasoning dataset (AbsR) together with a meaningful learning paradigm to teach LLMs how to leverage generic facts for reasoning purposes. The results show that our approach not only boosts the general reasoning performance of LLMs but also makes considerable strides towards their capacity for abstract reasoning, moving beyond simple memorization or imitation to a more nuanced understanding and application of generic facts.
@@ 1 Introduction
Humans proceed meaningful learning to induce common patterns or high-level abstractions to acquire abstract reasoning abilities Lake et al. (2017); Qiu et al. (2023). Such capabilities allow us to apply broad principles across diverse situations, demonstrating a deep understanding and versatile application of knowledge. As shown in Figure 1 (a), humans can deduce “rock dissolved” and “solution turns orange” when given “adding rock into hydrochloric acid” and “adding rust into sulphuric acid”, respectively. This stems from the established expertise in human mind (the generic fact “acid is corrosive”), and extends to applications in different scenarios. Of late, remarkable headways of LLMs have pushed AI much further towards human-like intelligence Achiam et al. (2023); Touvron et al. (2023). Interestingly, can LLMs act like humans to instinctively consider the generic fact for flexible applications in different scenarios?
Figure 1: The responses of (a) Humans and (b) LLMs when facing two questions supported by a genric fact.
LLMs Size Vanilla Accuracy AbsAcc
LLaMA-2 7B 51.67 24.82
@@
Table 1: The vanilla accuracy and abstract reasoning accuracy (AbsAcc) of LLaMA-2 Touvron et al. (2023) and Orca-2 Mitra et al. (2023) on e-CARE Du et al. (2022). We will formally define AbsAcc in Sec 2.1.
In our pilot investigation, we observed that LLMs appear to lack satisfying capabilities in abstract reasoning. As illustrated in Table 1, there is a notable discrepancy—exceeding 20%—between vanilla accuracy and abstract reasoning accuracy (AbsAcc). Despite the extensive pre-training of LLMs, which equips them with a vast repository of generic facts, they seem unable to leverage this information as flexibly as humans do (Figure 1 (b)).
In this paper, we aim to systematically and quantitatively investigate the abstract reasoning of LLMs. We first formally define the abstract reasoning metric. Then we carry out abstract reasoning and knowledge probing tasks for a preliminary study. Our investigation comprises two main components: the reasoning tasks can quantify the abstract reasoning capabilities of LLMs. The generic fact probing task can give a further analysis from the perspective of generic fact mastery. Hereafter, we improve abstract reasoning of LLMs based on the analysis of the preliminary study. Unlike existing chain-of-thought (CoT) methods Wei et al. (2022a); Kojima et al. (2022), which conduct reasoning with uncontrolled step-by-step explanation, we create a dataset AbsR tailored for abstract reasoning. AbsR not only includes generic facts but also their guided explanations, offering a coached approach to understand the process of reasoning with common patterns. Finally, to make LLMs subconsciously exploit the generic fact like humans, we design a simple but effective learning paradigm called meaningful learning to simulate the implicit knowledge learning process Du et al. (2019), enabling LLMs to implicitly learn and utilize generic facts for reasoning under the guidance of generic facts and the generic-fact-guided explanation.
Extensive experiments across multiple reasoning and natural language understanding benchmarks, demonstrate that our proposed paradigm for meaningful learning not only improves general reasoning of LLMs, but also excels in abstract reasoning. This distinction in performance is particularly pronounced in the realm of abstract reasoning, underscoring the unique efficacy of our approach in nurturing the higher-order thinking skills that are essential for sophisticated cognitive processing in artificial intelligence systems. Further analysis and ablation studies yield additional evidence supporting the effectiveness of our approach.
We summarize our contributions as follows:
  • We provide a systematic and quantitative analysis of abstract reasoning in current LLMs, and we develop an abstract reasoning dataset with generic-fact-guided explanation.
  • We propose a meaningful learning paradigm, which can stimulate the abstract reasoning capabilities of existing LLMs.
  • We achieve significant improvement in general reasoning and abstract reasoning on a wide range of benchmarks.
Figure 2: The computation of abstract reasoning metric.
@@ 2.1 Abstract Reasoning Metric
Abstract reasoning requires models to apply general patterns of high-level abstractions to different scenarios or questions. Different from general reasoning, which focuses on each single example, abstract reasoning takes the examples supported by a generic fact as a whole. We define abstract reasoning accuracy due to the lack of a proper metric.
@@
As shown in Figure 2, for hard constraint, akin to Qiu et al. (2023), we suppose an LLM $\mathcal{LM}$ has grasped the generic fact $r_{i}$ if and only if $\mathcal{LM}$ can correctly answer all examples in $S_{i}$. Hence, we define the abstract reasoning accuracy as the proportion of generic facts that $\mathcal{LM}$ has grasped:
$\texttt{AbsAcc}=\frac{\sum_{i}\mathbb{I}(\Phi(L,S_{i})=|S_{i}|)}{n},$ (1)
@@
2.2 Large Language Models
We offer an overview of the LLMs employed in our preliminary and upcoming experiments:
[…]
@@ 2.3 Experiment I: Abstract Reasoning
Examples Generic Facts
Original Filtered Orginal Filtered
21,314 13,877 13,045 5,608
Table 2: The statistics of original and filtered e-CARE.
To achieve this goal, we choose a multiple-choice question answering dataset e-CARE Du et al. (2022) for implementation. e-CARE is a large-scale explainable causal reasoning dataset with a generic fact in each example, and a generic fact can support more than one example. The full set (train, test, and dev) of e-CARE is selected for experiments. We filter out the examples that the corresponding generic facts support only one example. Table 2 shows the statistics of e-CARE.
We follow zero-shot setting to prompt the LLMs for answer selection from multiple choices, and the generic facts are not provided in the inputs (refer to Appendix A.1). Table 1 shows the overall results:
(1) There is a significant disparity (more than 20%) between the vanilla accuracy and AbsAcc across all LLMs. This proves two aspects: on the one hand, vanilla accuracy is not ideal for estimating abstract reasoning. On the other hand, although LLMs could answer questions accurately, they still cannot capture the common patterns or generic facts behind the questions to reach the answers.
(2) Larger LLMs outperform smaller ones due to their richer and more accurate knowledge. Orca-2, fine-tuned on LLaMA-2 with extensive reasoning examples, notably outperforms LLaMA-2, indicating that supervised fine-tuning enhances both reasoning and abstract reasoning abilities. However, even with large-scale and reasoning-specific post-training, the significant gap between vanilla accuracy and AbsAcc remains unbridged.
@@ 2.4 Experiment II: Generic Fact Probing
To delve into the abstract reasoning from the perspective of generic fact matery, we design a generic fact probing task to ask LLMs whether they possess the given generic fact (refer to Appendix A.2). Drawing inspiration from knowledge uncertainty estimation Amayuelas et al. (2023), we formulate this task as a yes or no problem.
LLMs LLaMA-2 Orca-2
7B 13B 7B 13B
Accuracy 62.44 98.86 26.97 12.78
Table 3: The overall results of the generic facts probing.
We also use e-CARE and the above LLMs to carry out experiments. The overall results are shown in Table 3. We could conclude as follows:
(1) Both 7B and 13B LLaMA-2 have substantial knowledge reserves, with LLaMA-2-13B being particularly knowing nearly all generic facts in e-CARE. This indicates the poor AbsAcc might come from a lack of ability to utilize knowledge.
(2) In contrast, Orca-2 knows limited generic facts, and the larger the model, the less it knows. This contrasts sharply with its strong reasoning abilities (vanilla accuracy). We suppose training Orca-2 is more like rote learning rather than meaningful learning, which implies Orca-2 tends to lose knowledge while acquiring reasoning.
Category Metrics LLaMA-2 Orca-2
7B 13B 7B 13B
Known AbsAcc 34.26 43.51 64.98 81.51
Unknown AbsAcc 17.19 41.84 49.00 58.38
To delve into the relation between generic facts and abstract reasoning, we categorize the examples in e-CARE according to whether LLMs know the generic facts. The results are represented in Table 4, from which we can draw the following conclusions:
@@
(2) Nonetheless, most of the results in “Unknown” category still have undeniable performance, suggesting that LLMs may engage in reasoning based on suspicious correlations or rely on memory recall. We can reasonably speculate that LLMs may consciously conduct memorization when answering reasoning-related questions.
In a conclusion, we can enhance the abstract reasoning abilities of LLMs in two ways: (1) Knowledge, injecting generic facts into the parametric memory of LLMs or offering LLMs the generic fact through prompts; (2) Reasoning, teaching LLMs how to utilize the generic facts to provide more precise responses to questions.
Figure 3: Two examples of AbsR guided by the generic fact about “more powerful computer”.
@@ 3 AbsR: Abstract Reasoning Dataset
To enhance the abstract reasoning abilities of LLMs through Knowledge and Reasoning ways, we need a dataset that can not only provide generic facts, but also teach LLMs how to exploit the generic fact to derive the correct answer in different scenarios (meaningful learning). Since there is no suitable dataset to assist us in achieving the goals, we create an Abstract Reasoning dataset (AbsR).
@@ 3.1 Generic Fact Collection
To obtain such a dataset described above, we need to collect the generic fact first. We choose GenericsKB Bhakthavatsalam et al. (2020) as the foundation of AbsR to supply generic facts. GenericsKB is a large-scale knowledge base with over 3.4 million sentences (e.g. Dogs bark) expressing general truths, each of which also includes the corresponding concept (e.g. Dog) and the confidence score (between 0 and 1). Finally, we sample a collection of 4,613 generic facts from GenericsKB (details can refer to Appendix B).
@@ 3.2 Dataset Construction
To construct the whole dataset, and inspired by previous works Wang et al. (2022a); Xu et al. (2023), we choose GPT-4 Achiam et al. (2023) as our data annotator. The API we used is gpt-4-1106-preview.
Specifically, for each sampled generic fact $r_{i}$, we would ask GPT-4 to create samples $S_{i}=\{s_{1}^{i},\cdots,s_{m_{i}}^{i}|1\leq m_{i}\leq 3\}$ according to $r_{i}$. Each sample $s^{i}_{j}$ contains the following terms: (1) a question $X^{i}_{j}$ with a few options, (2) a response $Y^{i}_{j}$ with an answer and an explanation guided by $r_{i}$. All terms form a triple $s^{i}_{j}=<r_{i},X^{i}_{j},Y^{i}_{j}>$ for each sample. The prompt we used for dataset creation can refer to Appendix C. An illustration of the created examples is shown in Figure 3.
Finally, without loss of generality, given the $j$-th sample $s^{i}_{j}$ of generic fact $r_{i}$, we can create two kinds of examples for our proposed meaningful learning. One is predicting $Y^{i}_{j}$ given $<r_{i},X^{i}_{j}>$ (K-example), while the other is predicting $Y^{i}_{j}$ given only $X^{i}_{j}$ (R-example). The examples are used to implicitly enhance the abstract reasoning abilities of LLMs through Knowledge and Reasoning ways. Table 5 shows the statistics of AbsR.
Examples Questions Generic Facts
Train 18,020 14,010 4,613
Test 200 200 104
Table 5: The statistics of AbsR.
@@ 4 Method: Meaningful Learning
To imitate humans’ instinctive use of general facts in reasoning, we develop a simple but effective learning paradigm called meaningful learning (MeanLearn). It can enhance the abstract reasoning abilities of LLMs in Knowledge and Reasoning ways. This is inspired by the implicit knowledge learning process, which uses hidden variables to learn event background knowledge Du et al. (2019). MeanLearn can make LLMs implicitly learn generic facts and solve problems under the guidance of generic facts.
Specifically, given an LLM $\mathcal{LM}$ with parameters $\theta$ and the input $x$, which constructs a conditional probability for the outputs $y$:
@@
Hereafter, on the one hand, we train $\mathcal{LM}$ to learn policies $p_{\theta}(Y_{t}|X,Y_{<t})$ and $q_{\theta}(Y_{t}|X,r,Y_{<t})$ together to push them closer, enabling $\mathcal{LM}$ implicitly learn the generac fact $r$. On the other hand, $\mathcal{LM}$ can learn how to apply $r$ by learning to generate the explanation in $Y$. Finally, we can obtain better abstract reasoning abilities when only given the question $X$, just like humans instinctively answer questions with general facts.

### v2

# Meaningful Learning: Enhancing Abstract Reasoning in Large Language Models via Generic Fact Guidance

@@ Abstract
Large language models (LLMs) have developed impressive performance and strong explainability across various reasoning scenarios, marking a significant stride towards mimicking human-like intelligence. Despite this, when tasked with several simple questions supported by a generic fact, LLMs often struggle to abstract and apply the generic fact to provide consistent and precise answers, revealing a deficiency in abstract reasoning abilities. This has sparked a vigorous debate about whether LLMs are genuinely reasoning or merely memorizing. In light of this, we design a preliminary study to quantify and delve into the abstract reasoning abilities of existing LLMs. Our findings reveal a substantial discrepancy between their general reasoning and abstract reasoning performances. To relieve this problem, we tailor an abstract reasoning dataset (AbsR) together with a meaningful learning paradigm to teach LLMs how to leverage generic facts for reasoning purposes. The results show that our approach not only boosts the general reasoning performance of LLMs but also makes considerable strides towards their capacity for abstract reasoning, moving beyond simple memorization or imitation to a more nuanced understanding and application of generic facts. The code is available at https://github.com/Waste-Wood/MeanLearn.
@@ 1 Introduction
Humans proceed with meaningful learning to induce common patterns or high-level abstractions to acquire abstract reasoning abilities [29, 37]. Such capabilities allow us to apply broad principles across diverse situations, demonstrating a deep understanding and versatile application of knowledge. As shown in Figure 1 (a), humans can deduce “rock dissolved” and “the skin suffers pain” when given “adding rock into hydrochloric acid” and “acid touches human skin”, respectively. This stems from the established expertise in the human mind (the generic fact “acid is corrosive”), and extends to applications in different scenarios. Of late, remarkable headways of LLMs have pushed AI much further towards human-like intelligence [2, 47]. Interestingly, can LLMs act like humans to instinctively consider the generic fact for flexible applications in different scenarios?
In our pilot investigation, we observed that LLMs appear to lack satisfying capabilities in abstract reasoning. As illustrated in Table 1, there is a notable discrepancy—exceeding 17%—between vanilla accuracy and abstract reasoning accuracy (AbsAcc) for all LLMs, while humans show only a small disparity. Despite the extensive pre-training of LLMs, which equips them with a vast repository of generic facts, they seem unable to leverage this information as flexibly as humans do (Figure 1 (b)).
In this paper, we aim to systematically and quantitatively investigate the abstract reasoning of LLMs. We first formally define the abstract reasoning metric. Then we carry out abstract reasoning and knowledge probing tasks for a preliminary study. Our investigation comprises two main components: the reasoning tasks can quantify the abstract reasoning capabilities of LLMs. The generic fact probing task can give a further analysis from the perspective of generic fact mastery. Hereafter, we improve abstract reasoning in LLMs based on the analysis of the preliminary study. Unlike existing chain-of-thought (CoT) methods [52, 27], which conduct reasoning with uncontrolled step-by-step explanation, we create a dataset AbsR tailored for abstract reasoning. AbsR not only includes generic facts but also their guided explanations, offering a coached approach to understand the process of reasoning with common patterns. Finally, to make LLMs subconsciously exploit the generic fact like humans, we design a simple but effective learning paradigm called meaningful learning (MeanLearn) to simulate the implicit knowledge learning process [13], enabling LLMs to implicitly learn and utilize generic facts for reasoning without requiring generic facts as input.
We evaluate MeanLearn on six out-of-domain (OOD) reasoning and language understanding benchmarks. Experimental results demonstrate that MeanLearn not only improves general reasoning of LLMs but also excels in abstract reasoning. This distinction in performance is particularly pronounced in the realm of abstract reasoning, underscoring the unique efficacy of our approach in nurturing the higher-order thinking skills that are essential for sophisticated cognitive processing in artificial intelligence systems. Further analysis and ablation studies yield additional evidence supporting the effectiveness of our approach. We summarize our contributions as follows:
  • We provide a systematic and quantitative analysis of abstract reasoning in current LLMs, and we develop an abstract reasoning dataset with generic-fact-guided explanations.
  • We propose meaningful learning to improve abstract reasoning in LLMs with generic fact.
  • We achieve significant improvement in general reasoning and abstract reasoning on various OOD reasoning and language understanding benchmarks.
Figure 1: The responses of (a) Humans and (b) LLMs when facing two questions which are supported by the same generic fact.
Table 1: The vanilla accuracy and abstract reasoning accuracy (AbsAcc) on e-CARE [14]. We will formally define AbsAcc in Sec. 2.1.
Methods Size Vanilla Accuracy AbsAcc
LLaMA-2 [47] 7B 51.67 24.82
@@
GPT-3.5 [7] >20B 83.25 66.08
Human [14] 92.00 89.90
Figure 2: Computation of abstract reasoning metric.
@@ 2.1 Abstract reasoning metric
Abstract reasoning requires models to apply general patterns or high-level abstractions to different scenarios or questions. Different from general reasoning, which focuses on each single example, abstract reasoning takes the examples supported by a generic fact as a whole. We define abstract reasoning accuracy due to the lack of a proper metric.
@@
As shown in Figure 2, akin to Qiu et al. [37], we suppose an LLM $\mathcal{LM}$ has grasped the generic fact $r_{i}$ if and only if $\mathcal{LM}$ can correctly answer all examples in $S_{i}$. Hence, we define abstract reasoning accuracy (denoted as AbsAcc) as the proportion of generic facts that $\mathcal{LM}$ has grasped:
$\texttt{AbsAcc}=\frac{\sum_{i}\mathbb{I}(\Phi(\mathcal{LM},S_{i})=|S_{i}|)}{n},$ (1)
@@
2.2 Large language models for evaluation
We employ several LLMs in our preliminary experiments: (1) LLaMA-2 [47] is an open-access LLM. We use 7B and 13B LLaMA-2 for experiments; (2) Orca-2 [35] is an open-access LLM finetuned on LLaMA-2 with over 80K reasoning-specific examples. We use 7B and 13B Orca-2 for experiments. (3) GPT-3.5 [7] is a limited access LLM, we use gpt-3.5-turbo-0125 for experiments.
Table 2: Accuracy of generic fact probing.
LLMs LLaMA-2 Orca-2 GPT-3.5
7B 13B 7B 13B >20B
Accu. 62.44 70.32 26.97 12.78 77.30
Category Metrics LLaMA-2 Orca-2 GPT-3.5
7B 13B 7B 13B >20B
Known AbsAcc 34.26 43.51 64.98 81.51 68.31
Unknown AbsAcc 17.19 31.84 49.00 58.38 58.49
@@ 2.3 Experiment I: abstract reasoning
To achieve this goal, we choose a multiple-choice question-answering dataset e-CARE [14] for implementation. e-CARE is a large-scale explainable causal reasoning dataset with a generic fact in each example, and a generic fact can support more than one example. The full set (train, test, and dev) of e-CARE is selected for experiments. We filter out the examples that the corresponding generic facts support only one example. Finally, we obtain 13,045 examples supported by 5,608 generic facts.
We follow zero-shot setting to prompt the LLMs for answer selection from multiple choices, and the generic facts are not provided in the inputs (refer to Appendix B.1). Table 1 shows the overall results:
(1) The significant disparity (over 17%) between the vanilla accuracy and AbsAcc for all LLMs proves two aspects: on the one hand, vanilla accuracy is not ideal for estimating abstract reasoning. On the other hand, although LLMs could answer questions accurately, they still cannot capture the common patterns or generic facts behind the questions.
(2) Larger LLMs outperform smaller ones due to their richer knowledge and stronger reasoning ability. Orca-2, additional trained on LLaMA-2, notably outperforms LLaMA-2, indicating that supervised fine-tuning enhances both reasoning and abstract reasoning abilities. However, even with large-scale post-training, the significant gap between vanilla accuracy and AbsAcc remains unbridged.
(3) The minimal gap between vanilla accuracy and AbsAcc for humans suggests the large gap for LLMs is not due to varying difficulties of the example supported by the same generic fact.
@@ 2.4 Experiment II: generic fact probing
To delve into the abstract reasoning from the perspective of generic fact mastery, we design a generic fact probing task to ask LLMs whether they possess the given generic fact (refer to Appendix B.2 for more details). Drawing inspiration from knowledge uncertainty estimation [3], we formulate this task as a yes or no problem. The overall results are shown in Table 3. We could conclude as follows:
(1) LLaMA-2 and GPT-3.5 have substantial knowledge reserves. This indicates the poor AbsAcc of LLaMA-2 might come from a lack of ability to apply the generic fact to different reasoning scenarios. This problem also exists in GPT-3.5, but it is not as serious as LLaMA-2.
(2) In contrast, Orca-2 knows limited generic facts, and the larger the model, the less it knows. This contrasts sharply with its strong reasoning abilities (vanilla accuracy). We suppose training Orca-2 is more like rote learning rather than meaningful learning, which implies that Orca-2 tends to lose knowledge while acquiring reasoning.
To deeply investigate the relation between generic facts and abstract reasoning, we categorize the examples in e-CARE according to whether LLMs know the generic facts. The results are represented in Table 3, from which we can draw the following conclusions:
@@
(2) Nonetheless, most of the results in the “Unknown” category still have undeniable performance, suggesting that LLMs may engage in reasoning based on suspicious correlations or rely on memorization.
In conclusion, we can enhance the abstract reasoning abilities of LLMs in two ways: (1) Knowledge, injecting generic facts through additional training or prompting; (2) Reasoning, teaching LLMs how to utilize the generic facts to provide more precise responses to questions.
Figure 3: Two samples of AbsR guided by the generic fact about “unusual conditions”.
@@ 3 AbsR: abstract reasoning dataset
To enhance the abstract reasoning abilities of LLMs through Knowledge and Reasoning ways, we need a dataset that can not only provide generic facts but also teach LLMs how to exploit the generic fact to derive the correct answer in different scenarios (meaningful learning). Since there is no suitable dataset to assist us in achieving the goals, we create an Abstract Reasoning dataset (AbsR).
@@ 3.1 Generic fact collection
To obtain the generic facts, We chose GenericsKB [5] as the foundational generic fact base for AbsR. GenericsKB is a large-scale knowledge base with over 3.4 million sentences (e.g. Dogs bark) expressing general truths, each of which also includes the corresponding concept (e.g. Dog) and the confidence score (between 0 and 1). Hence, we filter and then sample a collection of 4,613 high-quality generic facts of different concepts from GenericsKB (details of generic fact filtering and sampling can refer to Appendix C).
Table 4: The statistics of AbsR.
Examples Questions Generic Facts
Train 18,020 9,010 4,613
Test 200 200 104
@@ 3.2 Dataset construction
To construct the whole dataset, and inspired by previous works [49, 56], we choose GPT-4 [2] as our data annotator. The API we used is gpt-4-1106-preview.
Specifically, for each sampled generic fact $r_{i}$, we would ask GPT-4 to create samples $S_{i}=\{s_{1}^{i},\cdots,s_{m_{i}}^{i}|1\leq m_{i}\leq 3\}$ in various scenarios based on $r_{i}$. Each sample $s^{i}_{j}$ contains the following terms: (1) a question $X^{i}_{j}$ with a few options, (2) a response $Y^{i}_{j}$ with an answer and an explanation guided by $r_{i}$. All terms form a triple $s^{i}_{j}=<X^{i}_{j},r_{i},Y^{i}_{j}>$ for each sample. The prompt for dataset creation can refer to Appendix D. Figure 3 shows an illustration of the created samples.
Finally, without loss of generality, given the $j$-th sample $s^{i}_{j}$ of generic fact $r_{i}$, we can create two kinds of examples for meaningful learning. One is predicting $Y^{i}_{j}$ given $<X^{i}_{j},r_{i}>$ (K-example), while the other is predicting $Y^{i}_{j}$ given only $X^{i}_{j}$ (R-example). The examples can implicitly enhance abstract reasoning in LLMs through Knowledge and Reasoning ways. Table 4 shows the statistics of AbsR.
3.3 The quality of AbsR
We conduct human evaluations to measure the quality of AbsR from the following dimensions (details and more statistics such as evaluation criteria, agreements, and pay can refer to Appendix E and F):
  • Human Performance: humans can achieve vanilla accuracy of 95% and AbsAcc of 93.27%;
  • Support Rate: 89% (the rate of examples which can be supported by the generic fact);
  • Diversity: 88.5% (the diversity of the examples supported by the same generic fact. Greater diversity indicates greater differences among the examples of the same generic fact).
For comparison, e-CARE [14] is also human-annotated, with examples generated from generic facts. On e-CARE, humans reached 92% vanilla accuracy, 89.9% AbsAcc, and an 87% support rate [14], showing that AbsR matches the quality of the human-annotated dataset.
@@ 4 Method: meaningful learning
To imitate humans’ instinctive use of general facts in reasoning, we develop a simple but effective learning paradigm called meaningful learning (MeanLearn). It can enhance the abstract reasoning abilities of LLMs in Knowledge and Reasoning ways. This is inspired by the implicit knowledge learning process, which uses hidden variables to learn event background knowledge [13]. MeanLearn can make LLMs implicitly learn generic facts and solve problems under the guidance of generic facts.
Specifically, an LLM $\mathcal{LM}$ with parameters $\theta$ and the input $x$ can construct a conditional probability for the output $y$:
@@
Hereafter, on the one hand, we train $\mathcal{LM}$ to learn policies $p_{\theta}(Y_{t}|X,Y_{<t})$ and $q_{\theta}(Y_{t}|X,r,Y_{<t})$ together to enable $\mathcal{LM}$ implicitly learn the generac fact $r$. On the other hand, $\mathcal{LM}$ can learn how to apply $r$ into different scenarios by learning the explanations in $Y$ of different $X$ supported by $r$. Finally, we can reason with the implicit guidance of $r$ when only given $X$, just like humans instinctively answer questions without explicitly giving the general facts.